how many milliliters of pure water should a pharmacist add to 50 ml of a 15% minoxidil solution to make a 10% minoxidil solution?
Answer by stanbon(75887)

how many milliliters of pure water should a pharmacist add to 50 ml of a 15% minoxidil solution to make a 10% minoxidil solution?
-----
Equation:
water + water = water
------
0.85*50 + 1*x = 0.9*(50+x)
------
85*50 + 100x = 90*50 + 90x
-------
10x = 5*50
x = 25 ml (amt. of water to add)
--------------
Cheers,
Stan H.
----